# Extending class group action attacks via sesquilinear pairings

### v1

@@ 5 Recovering partial torsion image information
Let $E$ and $E^{\prime}$ be $\mathcal{O}$-oriented supersingular curves over $\overline{\mathbb{F}}_{p}$ upon which we can efficiently compute the action of a generator $\tau$ for $\mathcal{O}$. Assume that the discrete logarithm in $\mu_{m}$ and group operations on $E[m]$ are efficient. Assume also that $E[m]$ is a cyclic $\mathcal{O}$-module, and that the hidden oriented isogeny $\phi:E\rightarrow E^{\prime}$ has known degree coprime to $m$. Suppose we are given $P$ and $P^{\prime}$ such that $\mathcal{O}P=E[m]$ and $\mathcal{O}P^{\prime}=E^{\prime}[m]$. Then we can efficiently recover $N(\lambda)$ modulo $m$ for $\lambda\in\mathcal{O}$ such that $\phi P=[\lambda]P^{\prime}$.
@@
Note that by Theorem 4.3, $\widehat{T}^{\tau}_{m}(P,P)$ has order $m$. Since $m$ is smooth, using the reduced pairing, we can solve a discrete log problem in $\mu_{m}$ to obtain $N(\lambda)$ modulo $m$.
@@
This result would provide a square-root improvement on an exhaustive search over image points of $P$. More precisely, one could attack the class group action problem by trying all possible image points $\phi P$ for $P$, infer $\phi[\tau]P=[\tau]\phi P$, and therefore use the image of the $m$-torsion as input to the SIDH attacks, checking for success at each attempt. This is similar to [14, Section 4.1], for example. Here, the knowledge of $N(\lambda)$ restricts $\phi P$ to only $O(m)$ possible images, not $m^{2}$. This may provide another $O(\sqrt{p})$ runtime classical attack on CSIDH, for example. Attacks with this runtime already exist using random walks [15, 16]. To run such an exhaustive search SIDH attack, we need the degree of $\phi$ to be known and $m^{2}>\deg\phi$, $m$ to be coprime to $\deg\phi$, and $m$ to be smooth. Since we have great freedom in choosing $m$, we can expect to choose an $m$ around $\sqrt{\deg\phi}$. Then we have to try approximately $m$ different images $\lambda P^{\prime}$ of $P$ which give $N(\lambda)$ as computed, running the SIDH attack on each. This attack applies very generally in the setup of the theorem above, whereas the attacks of [16] require a variety of hypotheses.
@@
This and other similar results in this paper and in [13] are a caution against decision Diffie-Hellman problems in which one must decide if a given point is the image point of a specified torsion point under a hidden isogeny. A result like the previous one reduces the possibilities for the torsion image (without pinning it down entirely). For an example, the IND-CPA hardness of SiGamal [21] depends upon such a problem, called the P-CSSDDH assumption. This is discussed in [4, Section 6.1], where the authors lament the triviality of the available self-pairings. There are non-trivial pairings of the type $\widehat{T}$ and $\widehat{W}$ which would apply to the SiGamal situation, but only if we had access to a different orientation on the curves and isogeny. The Frobenius orientation used in the P-CSSDDH assumption results in a trivial pairing once again, because the torsion is contained in the base field.
@@
There is a sense in which we cannot hope to obtain more information than $N(\lambda)$ modulo $m$ using these methods. If we post-compose our isogeny with an endomorphism from $\mathcal{O}$ of norm $1$ modulo $m$, then we do not change the degree modulo $m$, but we do change $\lambda$, replacing it with another $\lambda^{\prime}$ having the same norm modulo $m$. To detect the difference, we must feed in more information than just the degree modulo $m$. In fact, it is possible to recover the same result by a different method. Take a basis for $E$ and $E^{\prime}$ and change basis so that the Weil pairing takes a canonical diagonal form. Then the set of possible endomorphisms in $\mathcal{O}$ that preserve this diagonal form turns out to be the same ‘degree of freedom’ of $\lambda$ observed above. The pairings from [4] can be seen as getting around this by assuming $\lambda\in\mathbb{Z}$, in which case $N(\lambda)$ pins down $\lambda$ more effectively.
@@
In principle, the result above doesn’t require using $\widehat{T}$; it could be phrased in terms of one of the coordinates in Theorem 4.1 (5). This wouldn’t violate the classification of cyclic self-pairings in [4] because the domain is not $\mathbb{Z}$-cyclic.
@@ 6 Reduction from SIDH1 to SIDH
In [13], the authors consider a variety of variants on the SIDH problem which can be parameterized by the information available about the action of $\phi$ on the $m$-torsion (level structure information). In particular, they define the following problem.
@@
In either case we refer to $m$ as the level of the SIDH or SIDH1 problem. The authors of [13] show that if $m$ has a large smooth square factor, then SIDH1 of level $m$ (a single torsion point image of order $m$) reduces to SIDH of level $O(\sqrt{m})$ (two torsion point images of order $O(\sqrt{m})$). More recently, a manuscript in preparation (presented at Caipi Symposium 2024 [3]) generalizes the SIDH attacks of [2, 20, 23], directly attacking SIDH1 without the requirement that $m$ have a large square factor. Both approaches require that $m>\deg\phi$.
@@
Let $E$ and $E^{\prime}$ be $\mathcal{O}$-oriented supersingular curves over $\overline{\mathbb{F}}_{p}$, upon which we can efficiently compute the action of endomorphisms from $\mathcal{O}$. Assume that $m$ is smooth and coprime to the discriminant. Assume also that $E[m]$ is a cyclic $\mathcal{O}$-module, and that the hidden isogeny $\phi:E\rightarrow E^{\prime}$ has known degree coprime to $m$. Then the problem SIDH1 of level $m$ to find $\phi$ reduces, in a polynomial number of operations in the field of definition of $E[m]$, to SIDH of level $m$ on the same curve $E$ and same $\phi$.
@@
We are given $\phi R$ for some point $R\in E[m]$ of order $m$. We wish to recover a second torsion point image, resulting in an SIDH problem. First, by Sunzi’s Theorem, we can reduce the problem to prime powers $m=p^{k}$. By assumption, $p$ is not ramified. Hence we may assume $p$ is a split or inert.
Case that $p$ is inert. We know $\mathcal{O}R$ is an $\mathcal{O}$-submodule of $E[m]\cong\mathcal{O}/m\mathcal{O}$. If $p$ is inert, it must be isomorphic to $\mathcal{O}/p^{s}\mathcal{O}$. However, $\mathcal{O}/p^{s}\mathcal{O}$ doesn’t have elements of additive order $p^{k}$ unless $s=k$. Thus $\mathcal{O}R=E[m]$. Given any other point $Q$, we may compute $\eta$ such that $Q=[\eta]R$ (using basis $R$ and $[\tau]R$). Then $\phi Q=\phi[\eta]R=[\eta]\phi R$.
Case that $p$ is split. Write $m=p^{k}=\mathfrak{b}\overline{\mathfrak{b}}$, where $N(\mathfrak{b})=m$. Then we have, without loss of generality, $S,T$ a basis for $E[m]$ so that $\mathfrak{b}T=\{\mathcal{O}\}$, and $\overline{\mathfrak{b}}S=\{\mathcal{O}\}$. Similarly, let $S^{\prime}$ and $T^{\prime}$ be a basis for $E[m]$ so that $\mathfrak{b}T^{\prime}=\{\mathcal{O}\}$ and $\overline{\mathfrak{b}}S^{\prime}=\{\mathcal{O}\}$. To find such a basis, one could use Remark 3 or more standard methods.
Now the mapping $\phi$, as a matrix from basis $S,T$ to basis $S^{\prime},T^{\prime}$, is diagonal, with some integers $k_{1}$ and $k_{2}$ on the diagonal. By writing $R$ and $\phi R$ in the relevant bases, namely $R=[a]S+[b]T$, $\phi R=[c]S^{\prime}+[d]T^{\prime}$, we learn that $ak_{1}\equiv c,bk_{2}\equiv d\pmod{m}$. We also know that $\deg\phi\equiv k_{1}k_{2}\pmod{m}$. Without loss of generality, at least one of $a$ or $b$ is coprime to $m$, so we know at least one of $k_{1}$ or $k_{2}$, and the degree equation then gives us the other.
@@ 7 Diagonal SIDH
The following problem arises in [13, Lemma 6 and Section 5.6].
@@
Fix $d,m\in\mathbb{Z}$. Let $E,E^{\prime}$ be elliptic curves defined over $\mathbb{F}_{q}$, where $m$ is coprime to $q$. Let $P,Q\in E[m]$ form a basis. Suppose there exists an isogeny $\phi:E\rightarrow E^{\prime}$ of known degree $d$. Suppose that generators $P^{\prime}$ of $\langle\phi P\rangle$ and $Q^{\prime}$ of $\langle\phi Q\rangle$ are known.
@@
Suppose $E$ and $E^{\prime}$ are $\mathcal{O}$-oriented (and one can compute the action of the endomorphisms efficiently, as usual). Let $m$ be a smooth integer such that modulo $m$, $1$ has polynomially many square roots. Then Diagonal SIDH with known degree for an oriented isogeny $\phi:E\rightarrow E^{\prime}$ is solvable in polynomial time, provided $\mathcal{O}P=E[m]$ or $\mathcal{O}Q=E[m]$.
@@
Let the Diagonal SIDH problem be given in terms of basis $P,Q$ for $E[m]$ and generators $P^{\prime}$ and $Q^{\prime}$ for $\langle\phi P\rangle$ and $\langle\phi Q\rangle$ respectively. Assume without loss of generality that $\mathcal{O}P=E[m]$. Then by Theorem 5.1, we can efficiently recover $N(\lambda)$ such that $\phi P=\lambda P^{\prime}$. However, the Diagonal SIDH setup guarantees that $\lambda\in\mathbb{Z}$, hence we have recovered $\lambda^{2}$ modulo $m$. By assumption, this gives only polynomially many possible values for $\lambda$, each of which can be tested by running the SIDH attacks, until one recovers $\phi$.
An instance of the Diagonal SIDH problem is the problem underlying the FESTA cryptosystem [1, Problem 7]. In this case $m$ is chosen to be a power of $2$, so the attack above would apply if FESTA were instantiated in a situation where the isogeny was oriented (for known orientations). Assuming an $\mathcal{O}$-orientation, the condition $\mathcal{O}P=E[m]$ or $\mathcal{O}Q=E[m]$ is reasonably likely to occur by chance if not explicitly avoided.
@@
The hard problem underlying FESTA reduces to finding explicit $\mathcal{O}$-orientations of the curves $E$ and $E^{\prime}$ and the isogeny $\phi$ between them.
@@
In [13, Section 5.6], it is shown how to reformulate the problem of finding an isogeny of fixed degree $d$ between oriented curves (the class group action problem) as a Diagonal SIDH problem, where $m$ is a product of primes split in $\mathcal{O}$. The method of reduction, in brief, uses the eigenspaces associated to a split prime in the orientation, which must map to each other. However, the conditions under which Theorem 7.1 applies – that $m$ have few square roots, and $P$ or $Q$ be generators of $E[m]$ as an $\mathcal{O}$-module – both fail in the Diagonal SIDH problems that result from the reduction of [13]. This means we cannot chain these attacks together to attack class group action problems!
8 When $N$ divides the discriminant
Suppose $N\mid\Delta_{\mathcal{O}}$, where $N=N(\tau)$ for $\tau\in\mathcal{O}$. In this case the pairing $\widehat{T}^{\tau}_{N}$ becomes trivial. However, a modification is more interesting. Let $n\in\mathbb{Z}$, and define
$T^{\prime}_{n}(P,Q)=\left(t_{n}([\tau]P,Q),t_{n}(P,Q)\right)\in((\mathbb{F}^{*})/(\mathbb{F}^{*})^{n}))^{\times 2}.$
This modification does not preserve all of the properties of Theorem 4.1 but importantly, it inherits compatibility from $t_{n}$, so that for $\phi:E\rightarrow E^{\prime}$ compatible with $\mathcal{O}$, we have
$T^{\prime}_{n}(\phi Q,\phi Q)=T^{\prime}_{n}(Q,Q)^{\deg\phi}.$
In [4], the authors use generalized pairings to determine the image of a single torsion point in $E[N]$, and then reduce to SIDH with $E[\sqrt{N}]$ when $N$ is a smooth square. In fact, recent further development of the SIDH attacks (in preparation [3]) generalize to image information on subgroups of a large enough size, not just full torsion subgroups, which effectively removes the restriction that $N$ be square.
Inspired by this result, we develop a similar reduction using the pairing above. The main advantage of our situation over that in [4] is the computation of the pairing, which requires only operations in the field of definition of $E[N]$. Because the pairings used in [4] may require a move to the field of definition of $E[N^{2}]$, our pairings result in a speedup in cases where that field of definition is large.
Lemma 2
Let $E$ have an orientation by $\mathcal{O}=\mathbb{Z}[\tau]$, and let $P\in E[m]$ be such that $\langle[\tau]P\rangle\cap\langle P\rangle=\{0\}$. Let $t$ be the minimal positive integer such that $[t]E[m]\subseteq\mathcal{O}P$. Let $f$ be the relative conductor of $\mathcal{O}$, so that $\tau=f\tau^{\prime}$ for another endomorphism $\tau^{\prime}$. Let $g=\gcd(m,f)$ and let $n=\gcd(N(\tau^{\prime}),m)$. Then, $g\leq t\leq gn$.
Proof
As a direct consequence of the hypotheses, $[\tau]P=[f][\tau^{\prime}]P\in E[m/g]$. On the other hand, $[\tau]P$ has order at least $m/gn$. Therefore $[\tau]P$ has order $s$ satisfying $m/gn\leq s\leq m/g$. By assumption, $\langle[\tau]P\rangle\cap\langle P\rangle=\{0\}$. Therefore, $s$ is the maximal integer dividing $m$ such that $E[s]\subseteq\mathcal{O}P$. But $E[s]=[m/s]E[m]$.

### v2

@@ 5 Recovering partial torsion image information
Let $E$ and $E^{\prime}$ be $\mathcal{O}$-oriented supersingular curves over $\overline{\mathbb{F}}_{p}$ upon which we can efficiently compute the action of a generator $\tau$ for $\mathcal{O}$. Assume that the discrete logarithm in $\mu_{m}$ is efficiently computable. Assume also that $E[m]$ is a cyclic $\mathcal{O}$-module, and that the hidden oriented isogeny $\phi:E\rightarrow E^{\prime}$ has known degree coprime to $m$. Suppose we are given $P$ and $P^{\prime}$ such that $\mathcal{O}P=E[m]$ and $\mathcal{O}P^{\prime}=E^{\prime}[m]$. Then we can efficiently recover $N(\lambda)$ modulo $m$ for $\lambda\in\mathcal{O}$ such that $\phi P=[\lambda]P^{\prime}$.
@@
Note that by Theorem 4.3, $\widehat{T}^{\tau}_{m}(P,P)$ has order $m$. Using the reduced pairing, we can solve a discrete log problem in $\mu_{m}$ to obtain $N(\lambda)$ modulo $m$. ∎
@@
This result improves upon a naïve exhaustive search over the possible images of a general point $P$ (on account of Theorem 4.3, we cannot use an eigenvector). More precisely, one could attack the class group action problem by trying all possible image points $\phi P$ for $P$, infer $\phi[\tau]P=[\tau]\phi P$, and use the imputed image of $E[m]$ for the SIDH attacks, checking for success at each attempt. This is similar to [FMP23, Section 4.1], for example. Here, the knowledge of $N(\lambda)$ restricts $\phi P$ to typically around $m$ possible images (between $m\prod_{\text{prime }q\mid m}(1-1/q)$ and $m\prod_{\text{prime }q\mid m}(1+1/q)$), rather than all $m^{2}$. To run such an attack, we need the degree of $\phi$ to be known and $m^{2}>\deg\phi$, $m$ to be coprime to $\deg\phi$, and $m$ to be smooth. Since we have great freedom in choosing $m$, we can expect to choose an $m$ around $\sqrt{\deg\phi}$.
The example of [CHM+23] described by (1) in Section 2.3 shows that when $m$ is a power of a prime $\ell$ that splits in $\mathbb{Q}(\sqrt{-p})$, the classical $m$-Weil pairing also provides an attack with this runtime. However, with the sesquilinear pairing, one does not require splitting conditions.
@@
This and other similar results in this paper and in [FFP24] are a caution against Decisional Diffie-Hellman problems in which one must decide if a given point is the image point of a specified torsion point under a hidden isogeny. A result like the previous one reduces the possibilities for the torsion image (without pinning it down entirely). For an example, the IND-CPA hardness of SiGamal [MOT20] depends upon such a problem, called the P-CSSDDH assumption. This is discussed in [CHM+23, Section 6.1], where the authors lament the triviality of the available self-pairings. There are non-trivial pairings of the type $\widehat{T}$ which would apply to the SiGamal situation, but only if we had access to a different orientation on the curves and isogeny. The Frobenius orientation used in the P-CSSDDH assumption results in a trivial pairing once again, because the torsion is contained in the base field.
@@
There is a sense in which we cannot hope to obtain more information than $N(\lambda)$ modulo $m$ using these methods. If we post-compose our isogeny with an endomorphism from $\mathcal{O}$ of norm $1$ modulo $m$, then we do not change the degree modulo $m$, but we do change $\lambda$, replacing it with another $\lambda^{\prime}$ having the same norm modulo $m$. To detect the difference, we must feed in more information than just the degree modulo $m$. In fact, it is possible to recover the same result by a different method. Take a basis for $E$ and $E^{\prime}$ and change basis so that the Weil pairing takes a canonical diagonal form. Then the set of possible endomorphisms in $\mathcal{O}$ that preserve this diagonal form turns out to be the same ‘degree of freedom’ of $\lambda$ observed above. The pairings from [CHM+23] can be seen as getting around this by assuming $\lambda\in\mathbb{Z}$, in which case $N(\lambda)$ pins down $\lambda$ more effectively.
@@
In principle, the result above doesn’t require using $\widehat{T}$; it could be phrased in terms of one of the coordinates in Theorem 4.1 (5). This wouldn’t violate the classification of cyclic self-pairings in [CHM+23] because the domain is not $\mathbb{Z}$-cyclic.
@@ 6 Reduction from SIDH1 to SIDH
In [FFP24], the authors consider a variety of variants on the SIDH problem which can be parameterized by level structure for on the $m$-torsion preserved by $\phi$. In particular, they define the following problem.
@@
In either case we refer to $m$ as the level of the SIDH or SIDH1 problem. The authors of [FFP24] show that if $m$ has a large smooth square factor, then SIDH1 of level $m$ (a single torsion point image of order $m$) reduces to SIDH of level $O(\sqrt{m})$ (two torsion point images of order $O(\sqrt{m})$). More recently, a manuscript in preparation (presented at Caipi Symposium 2024 [CDM+24]) generalizes the SIDH attacks of [CD23, MMP+23, Rob23b], directly attacking SIDH1 without the requirement that $m$ have a large square factor. Both approaches require that $m>\deg\phi$.
@@
Although the proof relies on taking an ‘imaginary quadratic viewpoint,’ it does not make use of the sesquilinear pairings.
@@
Let $E$ and $E^{\prime}$ be $\mathcal{O}$-oriented supersingular curves over $\overline{\mathbb{F}}_{p}$, upon which we can efficiently compute the action of endomorphisms from $\mathcal{O}$. Assume that $m$ is smooth and coprime to the discriminant. Assume also that $E[m]$ is a cyclic $\mathcal{O}$-module, and that the hidden isogeny $\phi:E\rightarrow E^{\prime}$ has known degree coprime to $m$ and is compatible with the $\mathcal{O}$-orientations. Then the problem SIDH1 of level $m$ to find $\phi$ reduces, in a polynomial number of operations in the field of definition of $E[m]$, to SIDH of level $m$ on the same curve $E$ and same $\phi$.
@@
For convenience, write $\mathcal{O}=\mathbb{Z}[\tau]$. We are given $\phi R$ for some point $R\in E[m]$ of order $m$. We wish to recover a second torsion point image resulting in an SIDH problem. First, by Sunzi’s Theorem, we can reduce the problem to prime powers $m=q^{k}$. By assumption, $q$ is not ramified. Hence we may assume $q$ is split or inert.
Case that $q$ is inert. We know $\mathcal{O}R$ is an $\mathcal{O}$-submodule of $E[m]\cong\mathcal{O}/m\mathcal{O}$. If $q$ is inert, it must be isomorphic to $\mathcal{O}/q^{s}\mathcal{O}$. However, $\mathcal{O}/q^{s}\mathcal{O}$ doesn’t have elements of additive order $q^{k}$ unless $s=k$. Thus $\mathcal{O}R=E[m]$. Given any other point $Q$, we may compute $\eta$ such that $Q=[\eta]R$ (using basis $R$ and $[\tau]R$). Then $\phi Q=\phi[\eta]R=[\eta]\phi R$.
Case that $q$ is split. Write $m=q^{k}=\mathfrak{b}\overline{\mathfrak{b}}$, where $N(\mathfrak{b})=m$. Write $\ker\mathfrak{b}:=\{P\in E[m]:\beta P=\mathcal{O}\text{ for all }\beta\in\mathfrak{b}\}$, and similarly for $\overline{\mathfrak{b}}$. Then these are distinct cyclic subgroups of order $m$. Thus there exists a basis $S,T$ for $E[m]$ so that $T\in\ker\mathfrak{b}$ and $S\in\ker\overline{\mathfrak{b}}$. Similarly, let $S^{\prime}$ and $T^{\prime}$ be a basis for $E^{\prime}[m]$ so that $T^{\prime}\in\ker\mathfrak{b}$ and $S^{\prime}\in\ker\overline{\mathfrak{b}}$. To find such subgroups, one can use linear algebra, as follows. The problem of finding $\ker\mathfrak{b}$ can be rephrased as solving for coefficients $a$ and $b$ for $T=aP+bQ$ in terms of a basis $P,Q$ for $E[m]$, subject to linear conditions determined by the action of $\mathfrak{b}$, which we can make explicit in terms of the known action of $\tau$. In addition, by adding a gcd condition on the coefficients, one can choose $T$ to be of full order $m$.
Now the mapping $\phi$, as a matrix from basis $S,T$ to basis $S^{\prime},T^{\prime}$, is diagonal, with some integers $k_{1}$ and $k_{2}$ on the diagonal (as $\phi$ respects the $\mathcal{O}$-orientation). By writing $R$ and $\phi R$ in the relevant bases, namely $R=[a]S+[b]T$, $\phi R=[c]S^{\prime}+[d]T^{\prime}$, we learn that $ak_{1}\equiv c,bk_{2}\equiv d\pmod{m}$, where $a,b,c,d$ are known. We also know that $\deg\phi\equiv k_{1}k_{2}\pmod{m}$. Without loss of generality, at least one of $a$ or $b$ is coprime to $m$, so we know at least one of $k_{1}$ or $k_{2}$, and the degree equation then gives us the other. ∎
@@ 7 Diagonal SIDH
The following problem arises in [FFP24, Lemma 6 and Section 5.6].
@@
Fix $d,m\in\mathbb{Z}$. Let $E,E^{\prime}$ be elliptic curves defined over $\mathbb{F}_{q}$, where $m$ is coprime to $q$. Let $P,Q\in E[m]$ form a basis. Suppose there exists an isogeny $\phi:E\rightarrow E^{\prime}$ of known degree $d$. Suppose that generators $P^{\prime}$ of $\langle\phi P\rangle$ and $Q^{\prime}$ of $\langle\phi Q\rangle$ are known. Find $\phi$.
@@
Suppose $E$ and $E^{\prime}$ are $\mathcal{O}$-oriented (and one can compute the action of the endomorphisms efficiently, as usual). Let $m>4\deg\phi$ be a smooth integer such that modulo $m$, $1$ has polynomially many square roots. Then Diagonal SIDH with known degree for an oriented isogeny $\phi:E\rightarrow E^{\prime}$ is solvable in polynomial time, provided $\mathcal{O}P=E[m]$ or $\mathcal{O}Q=E[m]$.
@@
Let the Diagonal SIDH problem be given in terms of basis $P,Q$ for $E[m]$ and generators $P^{\prime}$ and $Q^{\prime}$ for $\langle\phi P\rangle$ and $\langle\phi Q\rangle$ respectively. Assume without loss of generality that $\mathcal{O}P=E[m]$. Then by Theorem 5.1, we can efficiently recover $N(\lambda)$ modulo $m$ such that $\phi P=[\lambda]P^{\prime}$. However, the Diagonal SIDH setup guarantees that $\lambda\in\mathbb{Z}$, hence we have recovered $\lambda^{2}$ modulo $m$. By assumption, this gives only polynomially many possible values for $\lambda$, each of which can be tested by running the SIDH attacks, until one recovers $\phi$.∎
An instance of the Diagonal SIDH problem is the problem underlying the FESTA cryptosystem [BMP23, Problem 7]. In this case $m$ is chosen to be a power of $2$, so the attack above would apply if FESTA were instantiated in a situation where the isogeny was oriented (for known orientations). Assuming an $\mathcal{O}$-orientation, the condition $\mathcal{O}P=E[m]$ or $\mathcal{O}Q=E[m]$ is reasonably likely to occur by chance if not explicitly avoided.
@@
The hard problem underlying FESTA, namely CIST (see [BMP23]), with $m>4\deg\phi$, reduces to finding explicit $\mathcal{O}$-orientations of the curves $E$ and $E^{\prime}$ respected by the isogeny $\phi$.
@@
In [FFP24, Section 5.6], it is shown how to reformulate the problem of finding an isogeny of fixed degree $d$ between oriented curves (the class group action problem) as a Diagonal SIDH problem, where $m$ is a product of primes split in $\mathcal{O}$. The method of reduction, in brief, uses the eigenspaces associated to a split prime in the orientation, which must map to each other. However, the conditions under which Theorem 7.1 applies – that $m$ have few square roots, and $P$ or $Q$ be generators of $E[m]$ as an $\mathcal{O}$-module – both fail in the Diagonal SIDH problems that result from the reduction of [FFP24]. This means we cannot chain these attacks together to attack class group action problems!
8 When $m$ divides the discriminant
Suppose $m\mid\Delta_{\mathcal{O}}$, where $m=N(\tau)$ for $\tau\in\mathcal{O}$. In this case the pairing $\widehat{T}^{\tau}_{m}$ becomes trivial. However, a modification is more interesting. Let $m\in\mathbb{Z}$, and define
$T^{\prime}_{m}:E[m](\mathbb{F})\times$ $E(\mathbb{F})/[m]E(\mathbb{F})\rightarrow((\mathbb{F}^{*})/(\mathbb{F}^{*})^{m})^{\times 2},$
$T^{\prime}_{m}(P,Q)$ $=\left(t_{m}([\tau]P,Q),t_{m}(P,Q)\right).$
This modification does not preserve all of the properties of Theorem 4.1 but importantly, it is bilinear and inherits compatibility from $t_{m}$, so that for $\phi:E\rightarrow E^{\prime}$ compatible with $\mathcal{O}$, we have
$T^{\prime}_{m}(\phi Q,\phi Q)=T^{\prime}_{m}(Q,Q)^{\deg\phi}.$
In [CHM+23], the authors use generalized pairings to determine the image of a single torsion point in $E[m]$, and then reduce to SIDH with $E[\sqrt{m}]$ when $m$ is a smooth square. As previously mentioned, recent further development of the SIDH attacks (in preparation [CDM+24]) generalize to image information on subgroups of a large enough size, not just full torsion subgroups, which effectively removes the restriction that $m$ be square.
Inspired by this result, we develop a similar reduction using the pairing above. The main advantage of our situation over that in [CHM+23] is the computation of the pairing, which requires only operations in the field of definition of $E[m]$. Because the pairings used in [CHM+23] may require a move to the field of definition of $E[m^{2}]$, our pairings result in a speedup in cases where that field of definition is large.